OpenStudy (ccastudentgrade10):

FANANDMEDAL let f(x)=x-2 and g(x)=x^2-3x+2 find f(x)*g(x) and find f(x)-g(x) i have to do a bunch of these and i dont understand how. my teacher didn't give us an example problem and she's out of office today.

OpenStudy (ccastudentgrade10):

@xMissAlyCatx

OpenStudy (xmissalycatx):

@pooja195 @imqwerty

OpenStudy (xmissalycatx):

Just plug in the f(x) and g(x) into the forms that they are given ...

OpenStudy (xmissalycatx):

And then simplify.

OpenStudy (ccastudentgrade10):

but how do i multiply x-2 and x^2-3x+2 ?

OpenStudy (ccastudentgrade10):

or subtract them

OpenStudy (xmissalycatx):

http://www.purplemath.com/modules/fcncomp3.htm Here, let this better explain it to you .

OpenStudy (xmissalycatx):

It should give you some proper information on how to do it..

OpenStudy (ccastudentgrade10):

i still dont understand, is there somewhere where they just do a problem thats the same to show you how to do it

OpenStudy (ccastudentgrade10):

like i just need once to be showed how to d it and then ill repeat it

OpenStudy (mathmale):

let f(x)=x-2 and g(x)=x^2-3x+2. Then (f*g)(x)= (x-2)*(x^2-3x+2). Try doing this; it's very similar to the "FOIL method" of multiplying binomials. (f-g)(x) is simple subtraction of like terms. f(x) = x - 2 -g(x) = -x^2 + 3x - 2 --------------------

OpenStudy (mathmale):

Please share your work.

OpenStudy (ccastudentgrade10):

would the subtraction one be -x^2 - 2x

OpenStudy (ccastudentgrade10):

@mathmale

OpenStudy (ccastudentgrade10):

for the multiplication one i just dont know where to start

OpenStudy (xmissalycatx):

Thank you @mathmale Sorry that I could not better explain it to ou @ccastudentgrade10

OpenStudy (ccastudentgrade10):

@mathmale